OpenStudy (yoloshroom):

2x^2-15x-8 how do i put this is standard form (Vertex Form)?

OpenStudy (lochana):

http://mathbitsnotebook.com/Algebra1/Quadratics/QDVertexForm.html read method 1 :)

OpenStudy (yoloshroom):

But i cant factor out the 2 so how do i do it?

OpenStudy (lochana):

say y = 2x^2-15x-8 y = 2(x^2 - 15x/2) - 8 y = 2(x^2 - 15x/2 + 225/16) -8 - 2(225/16) y = 2(x - 15/4)^2 - (8*16 + 450)/16 y = 2(x - 15/4)^2 - 289/8

OpenStudy (lochana):

please try to solve it on your own. you can't learn them by reading. You have to do it.:)

OpenStudy (yoloshroom):

I messed up the division.. (-15/2)/2

OpenStudy (lochana):

@YoloShroom have you found it?